# Simple crowd dynamics to generate complex temporal contact networks

### v1

@@ Abstract
Empirical contact networks or interaction networks demonstrate peculiar characteristics stemming from the fundamental social, psychological, physical mechanisms governing human interactions. Although these mechanisms are complex, we test whether we are able to reproduce some dynamical properties of these empirical networks from relatively simple models. In this study, we perform simulations for a range of 2D models of particle dynamics, namely the Random Walk, Active Brownian Particles, and Vicsek models, to generate artificial contact networks. We investigate temporal properties of these contact networks: the distributions of contact durations, inter-contact durations and number of contact per pair of particle. We demonstrate that the distribution of inter-contact durations can be recovered by the dynamics of these simple crowd particle models, and show that it is simply related to the well-know first-return process, which explains the $-3/2$ exponent that is found in both the numerical models and empirical contact networks.
@@ I Introduction
The structures and properties of temporal networks constrain processes which unfold on top of them. While understanding the effect of these structures is usually done with randomisation techniques [16, 18, 19, 14], generative models are needed to understand their origin. In the realm of generative models for temporal networks, there exist several noteworthy approaches, including but not limited to the following approaches. Holme introduced a fundamental technique known as “Static networks with link dynamics” [17, 39]. In this approach a temporal network is created by firstly generating a static network using a specific model and then generating a sequence of contacts for each link, typically without considering the network position of the links. Perra et al. [37] introduced an even more simplified model of temporal networks using a graph sequence framework. The approach involves increasing a time counter to $t$ and initializing an empty simple graph $G_{t}$. For each node $i$, activation occurs with a probability $a_{i}\Delta t$, connecting $i$ to $m$ other randomly selected nodes, whether active or not. The probability $a_{i}$ follows a truncated power-law distribution. The networks constructed in this way are called Activity-driven networks [28]. Other approaches tried to take into account the intrinsic heterogeneous properties of the dynamics of empirical networks by implementing reinforcement processes, such as link-node memory models [49], and Self-exciting point processes [31, 7].
@@ II.1 Models
Our primary objective is to understand the properties of empirical contact networks by exploring what happens for basic interaction models. One question in particular we aim to address is whether there exist universal properties that remain independent while varying the contact dynamics. To achieve this, we use three distinctive two-dimensional particle dynamics models: Random Walk, Active Brownian Particle, and Vicsek model. Fig. 1 shows a schematic explanation for the dynamics of a single particle between two consecutive time steps in each model. In all models we first assume that the particles are point-wise entities, which means their presence does not influence each other; their trajectories remain unchanged when they meet. In a second version, particles have a fixed size, and simply stop at contact.
@@
II.1.1 Random walk
In the 2D Random Walk, each walker selects a random direction at each subsequent step, uniformly distributed between 00 and $2\pi$, with a step length of $\Delta r$. The value of $\Delta r$ is drawn at random in the positive half of a Gaussian distribution of mean 00 and of variance $D*dt$, where $D$ and $dt$ represent the diffusion coefficient and time increment at each step respectively. This continuous randomness in the particle’s movements is classically known to be a good model for a brownian motion [27]. The diffusion coefficient allows us to control the characteristics of the random walk. For a given time increment $dt$, a larger diffusion coefficient implies that the random walker is more likely to move and spread rapidly throughout its surrounding space.
II.1.2 Active Brownian Particles (ABP)
The Active Brownian Particle (ABP) model adds a layer of complexity by adding a memory aspect into the particle motion. In the classical version of this model, particles undergo both a Brownian motion and a deliberate movement (usually the consequence of internal energy sources or external driving forces). In our simplified version, we exclude Brownian motion affecting position and focus solely on the self-propulsion of particles [40].
In practice, particles are constrained at each time step to select a new orientation uniformly within a specific range of angles, namely $\Delta\theta=[-\frac{\theta}{2},\frac{\theta}{2}]$ around their current orientation, while maintaining a constant velocity $v$.
In this model, trajectories of particles exhibit a persistence of the angle over the persistence length, which increases as the angle range $\Delta\theta$ decreases. Note that for $\Delta\theta=2\pi$, the ABP model is simply the 2D Random Walk with fixed step length.
II.1.3 Vicsek model
The Vicsek model focuses on collective behavior emerging from local interactions among self-propelled particles. Inspired by the movement of flocks of birds or schools of fish, this model assumes that particles align their velocities with those of their neighbors within a certain interaction radius $R_{n}$, while accounting for uncertainties caused by noise $\eta_{i}$. This alignment is mathematically described by the following equation:
$\Theta_{i}(t+\Delta t)=\langle\Theta_{j}\rangle_{|\mathbf{r_{i}}-\mathbf{r_{j}}|<R_{n}}+\eta_{i},$
where $\Theta_{i}(t)$ indicates the angle defining the direction of the velocity of particle $i$, while $\langle\Theta_{j}\rangle_{|\mathbf{r_{i}}-\mathbf{r_{j}}|<R_{n}}$ signifies the average angle of velocity vectors among the neighboring particles of particle $i$ within a distance of $R_{n}$. Hence, the particle moves at constant speed $v$ in new direction according to the following equation:
$\mathbf{r_{i}}(t+\Delta t)=\mathbf{r_{i}}(t)+\mathbf{v}\Delta t\binom{\cos\Theta_{i}(t)}{\sin\Theta_{i}(t)},$
where $\mathbf{r_{i}}(t)$ indicates the position of particle $i$ at time $t$.
Figure 1: Schematic representation of the motion mechanisms for the three models. (a) 2D Random Walk: at each time step, each particle selects a direction and a step length at random, following a uniform distribution for the angle and a half-gaussian for the step length. (b) Active Brownian Particle: at each time step, each particle updates its direction $\alpha$ by selecting uniformly at random a variation $\Delta\alpha$, with $\Delta\alpha$ constrained within the range $[-\frac{\theta}{2},\frac{\theta}{2}]$, while maintaining of a constant velocity $v$. The smaller $\theta$, the more the particle exhibits persistence of the direction of its motion. (c) Vicsek model: at each time step, particles align their direction with their neighbors by selecting the average direction over all neighbors within a given radius $R_{n}$. The alignment is altered by a noise $\eta$ similar to the ABP model.
II.2 Contact network
Figure 2: Typical four situations for particle contacts. When two particles A and B are closer than the detection radius $r_{D}$, four different situations can occur depending on their respective orientation vectors $D_{A}$ and $D_{B}$. As per our defined concept of contact, particles A and B are deemed to be in contact in configurations (a) and (b), as their are in both cases facing each other, i.e. in their respective front half-discs. In configurations (c) and (d) however, they are not in such a configuration and thus not considered to be in contact.
II.2.1 Empirical data
@@
II.2.2 Particle models
In order to be able to compare empirical and artificial interactions, we detect contacts in the models in a way that mirrors the contact definition from the SocioPatterns data. An interaction between particles is thus defined by two factors: the closeness of particles to each other and the orientation with which particles can detect each other. As illustrated in Fig. 2, particles can only detect another particle present in their front half circle, which is characterized by a detection radius $r_{D}$ and a detection vector named $\overrightarrow{D_{i}}$ for particle with index $i$. For the sake of simplicity, we assume that the detection vector for each particle coincides with the velocity vector of the particle. This assumption is logically sound, as in most contact cases the orientation of the line of sight coincides with the movement. We further assume a similar detection radius of $r_{D}=1$ for all particles.
Let us consider two particles $A$ and $B$ separated by a distance less than $r_{D}$. Let $\overrightarrow{AB}$ be the vector from particle $A$ to particle $B$. Particle B is then within the detection area of Particle A if $\overrightarrow{AB}\cdot\overrightarrow{D_{A}}\geq 0$. Similarly, Particle A is within the detection area of Particle B if $\overrightarrow{AB}\cdot\overrightarrow{D_{B}}\leq 0$. The fulfillment of both these conditions is necessary to establish that the two particles are in contact.
II.2.3 Contact networks comparison
To compare empirical and artificial contact networks, we focus on the distributions of three temporal properties of the contacts, namely contact duration, inter-contact duration and the number of contacts per pair of particles. In practice, for each pair of particles with indices $i$ and $j$, we monitor their contact occurrences over time, generating a timeline of their interactions. The timeline of contacts between particle $i$ and particle $j$ takes the following form: $[t_{1},t_{2},t_{3},t_{4},...,t_{k},...]$, where each odd-indexed element represents the beginning time of an interaction, and each even-indexed element represents the end time of the respective interaction. From this timeline array we can derive all the properties for a particular pair of particles, then join the results to obtain them for the whole population. For networks generated by the models, if the size of timeline array is an odd number (which indicates the contact was still ongoing at the end of the simulation), we simply drop the last element of the array.
@@ III Result
III.1 Temporal distributions form simple dynamics
@@ III.1.2 Inter-contact duration
However, we do observe something unexpected when examining the distributions of inter-contact duration. In all three models, this distribution exhibits the same power-law shape as it does in the data sets, and always with a similar exponent of $-1.5$ which is also observed for empirical contacts. For each model, this shape holds for all values of the parameter. The exponent is also the same in most cases, changing only slightly in the case of high diffusivity for 2DRW or long persistence for ABP. This remarkable fact seems to indicate that this particular property of contact dynamics between moving particles is largely independent from the details of the dynamics.
Focusing on the 2D random walk, arguments allow to relate the analytical derivation of this distribution to the distribution of first-passage times in a 1D random walk. Indeed, as they do not interact all particles have independent trajectories. The system can thus be reduced to a single pair of particles, for which we investigate the probability for two particles in contact to be in contact again after a time $\Delta\tau$. Classically, this 2D problem can be reduced to a 1D one by considering only the distance between the two particles. The evolution of this distance can be rewritten as a 1D random walk, which derivation gives a first passage time distribution with a power law tail of exponent $-1.5$. However, this method is applicable only for particles moving on an infinite space. In our case, the hard boundaries make it more difficult to use a similar approach. Nonetheless, the fact that we do get the same exponent for the tail of the distribution seems to indicate that the effect of the boundaries probably does not change the core of the derivation.
@@
Furthermore, the fact that we get the same exponent in the other models lets us formulate a general hypothesis about the universality of this effect: if the motion of particles is “random enough” to be statistically equivalent to a random walk, then the distribution of inter-contact duration exhibits a power-law tail with exponent $-1.5$.
@@
Finally, the fact that we observe the same feature in empirical data of contacts between individuals lets us believe that the motion of human beings, although far from a random walk at the individual’s scale and at short time, is also sufficiently “random” to exhibit this property as well when considering a crowd over a certain time window.
III.2 Effect of particle size and boundary conditions on inter-contact durations
Figure 4: Comparison of the distributions of inter-contact durations for the different version of the 2D Random Walk. Top: point particles, bottom: finite size particles. (a) Hard boundaries, (b) Periodic boundaries, (c) No boundaries.
In order to better understand the possible mechanisms responsible for the peculiar shape of the distribution of inter-contact durations, we focus on the 2D Random Walk and modified the way particles interact as well as the boundary conditions. For the particle size, the main difference between point particles and finite-size particles is the spatial organization: point particles can be as numerous as possible in a defined area of the simulated space and thus have unlimited simultaneous interactions, while the volume occupied by finite-size particles leads to a spatial organization and limits the number of simultaneous interactions. Changes in boundary conditions can modify the behaviour of a system. Hard boundaries impact the trajectories of particles while periodic boundaries do not; hard boundaries and periodic boundaries may impose finite-size effect on the global dynamics of the system while open boundaries do not.
As seen on Figure 4, changes in the particle size or in the boundary conditions do not affect at all the distribution of inter-contact durations, which retains in all cases its power law tail with an exponent $-1.5$. In particular, it works for the case of point particles doing a random walk with no boundaries. This very simple setting allows us to analytically derive the distribution.
III.3 Analytical argument
For the 2D Randow Walk of point particles on an infinite space, we first acknowledge that the trajectories of particles are uncorrelated, as they do not interact. The distribution of inter-contact durations for a population of particles is thus strictly identical as the one for only 2 particles.
Then, the system can be reduced to tracking the evolution of the distance $d$ between the two particles. If we define the random walk with the following process:
$\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{r}$ (1)
with $\mathbf{r}$ a random vector with an angle $\theta$ chosen uniformly in $[0,2\pi[$ and a norm $r$ chosen following the positive half of a gaussian, we have then:
$\mathbf{d}_{12}(t+1)=\mathbf{d}_{12}(t)+\mathbf{s}$ (2)
with $\mathbf{s}=\mathbf{r}_{2}-\mathbf{r_{1}}$. Given Eq. (1), $\mathbf{s}$ is also a random vector, which angle follows a uniform distribution. This indicates that $\mathbf{d}$ follows a 2D random walk process. As a consequence, the evolution of its norm $d$ can be written as a 1D random walk process.
The distribution of inter-contact durations is then the distribution of the time it takes for $d$, starting from $d\leq r_{D}$, to become smaller than the detection radius $r_{D}$ again. This distribution is exactly the distribution of first return times for the associated 1D random walk. Independently from the details of the stochastic process governing the evolution of $d$, this distribution has a power law tail $p(\Delta\tau)\sim\Delta\tau^{-3/2}$ [38], which is precisely the exponent we observe numerically.
@@ IV Conclusion
Acknowledgement
@@
Author contributions
M.G. proposed the core idea of the project. R.M., J.G. and M.G. contributed to the scientific discussions of the work. R.M. and J.G. wrote code and analyzed numerical and empirical data. R.M., J.G. and M.G. wrote and reviewed the final manuscript.
Competing interest

### v2

@@ Abstract
Empirical contact networks exhibit peculiar characteristics stemming from social, psychological, physical mechanisms governing human interactions. In this study, we test whether we are able to reproduce some of their dynamical properties from relatively simple models of 2D particle dynamics (random walk, active brownian particles and Vicsek model). Among the temporal properties of contact networks, we show that the distribution of inter-contact durations can be recovered with these simple models, and that this property is simply related to the well-know first-return process, which explains the $-3/2$ exponent that is found in both the numerical models and empirical contact networks.
@@ I Introduction
The structures and properties of temporal networks constrain processes which unfold on top of them. While understanding the effect of these structures is usually done with randomisation techniques [16, 18, 19, 14], generative models are needed to understand their origin. In the realm of generative models for temporal networks, there exist several noteworthy approaches, including but not limited to the following approaches. Holme introduced a fundamental technique known as “Static networks with link dynamics” [17, 39]. In this approach a temporal network is created by firstly generating a static network using a specific model and then generating a sequence of contacts for each link, typically without considering the network position of the links. Perra et al. [36] introduced an even more simplified model of temporal networks using a graph sequence framework, called activity-driven networks [27]. Other approaches tried to take into account the intrinsic heterogeneous properties of the dynamics of empirical networks by implementing reinforcement processes, such as link-node memory models [48], and Self-exciting point processes [30, 7].
@@ II.1 Models
Our primary objective is to understand the properties of empirical contact networks by exploring what happens for basic interaction models. One question in particular we aim to address is whether there exist universal properties that remain independent while varying the contact dynamics. To achieve this, we use three distinctive two-dimensional particle dynamics models: Random Walk, Active Brownian Particle, and Vicsek model. In all models we first assume that the particles are point-wise entities, which means their presence does not influence each other; their trajectories remain unchanged when they meet. In a second version, particles have a fixed size, and simply stop at contact.
@@
II.2 Contact networks
@@
In order to be able to compare empirical and artificial interactions, we detect contacts in the models in a way that mirrors the contact definition from the SocioPatterns data: two particles are in contact when they are closer than the detection radius and facing each other (see SI for details).
To compare empirical and artificial contact networks, we focus on the distributions of three temporal properties of the contacts, namely contact duration (number of consecutive time steps during which two particles are in contact), inter-contact duration (number of consecutive time steps during which two particles are not in contact) and the number of contacts per pair of particles.
@@ III Result
III.1 Temporal distributions for simple dynamics
@@ III.1.2 Inter-contact duration
However, we do observe something unexpected when examining the distributions of inter-contact duration. In all three models, this distribution exhibits the same power-law shape as it does in the data sets, and always with a similar exponent of $-1.5$ which is also observed for empirical contacts. For each model, this shape holds for all values of the parameter. The exponent is also the same in most cases, changing only slightly in the case of high diffusivity for 2DRW or long persistence for ABP. This remarkable fact seems to indicate that this particular property of contact dynamics between moving particles is largely independent from the details of the dynamics. Changing from point-particles to particles with a size which stop at contact, or modifying the boundary conditions to bi-periodic or infinite space does not modify this property (see SI).
Focusing on the random walk of point-particles on a boundless space, arguments allow to relate the analytical derivation of this distribution to the distribution of first-passage times in a 1D process describing the evolution of the distance between a pair of particles. In the Brownian limit, this distance approximates a one-dimensional diffusive process over short and long timescales [37] for which the first-passage time distribution is characterized by a power-law tail of exponent -3/2 [38].
@@
The fact that we numerically get the same exponent for the tail of the distribution with all boudary conditions and particle with or without sizes seems to indicate that the effect of these modifications does not change the core of the derivation. Furthermore, the fact that we get the same exponent in the other models (ABP and Vicsek before the phase transition) lets us formulate a general hypothesis about the universality of this effect: if the motion of particles is “random enough” to be statistically equivalent to a random walk, then the distribution of inter-contact duration exhibits a power-law tail with exponent $-3/2$.
@@
Finally, the fact that we observe the same feature in empirical data of contacts between individuals lets us believe that the motion of human beings, although far from being a random walk at the individual’s scale and at short time, is also “sufficiently random” to exhibit this property when considering a crowd over a certain time window.
@@ IV Conclusion
Acknowledgements
Author contributions
M.G. proposed the core idea of the project. R.M., J.G. and M.G. contributed to the scientific discussions of the work. R.M. and J.G. wrote code, performed numerical simulations and analyzed numerical and empirical data. R.M., J.G. and M.G. wrote and reviewed the final manuscript.
Funding
@@ Acknowledgements
Competing interests